# Optimal Kron-based Reduction of Networks (Opti-KRON) for Three-phase Distribution Feeders

### v1

@@ VI Simulation and Results
To evaluate the performance of the proposed framework, simulations are carried out on two realistic three-phase unbalanced distribution feeders from Vermont, USA: feeder A with 8,381 nodes and feeder B with 5,591 nodes. To investigate performance across different network scales, multiple sub-feeders of varying sizes (100, 300, 500, and 1000 nodes) are extracted from these systems. For each network, two representative operating scenarios are considered corresponding to the highest and lowest loading conditions from a week of historical hourly data, i.e. 168 loading scenarios. All computations were performed on the Vermont Advanced Computing Center (VACC) at the University of Vermont. The simulation codes used in this study were implemented in Julia 1.11.5. To solve the optimization problems, JuMP v1.29.1 [12] and the Gurobi Optimizer 12.0 were employed, with the optimality gap set to 0%. The GPU implementation was carried out using CUDA package [13].
@@
Across each test feeders, the achieved reduction levels and voltage errors ($0.001$ p.u.) are almost identical, which emphasize that increasing $q$ provides no meaningful gain in accuracy or reduction capability. On the other hand, small variations in SMICE suggest that larger $q$ values may slightly increase intra-cluster error due to reduced exploration of feasible assignments in each iteration. The remainder of the analysis, therefore focuses exclusively on $q=1$, which provides minimal computational burden.
Figure 5 compares the MILP-based and exhaustive search formulations on the 100-node sub-feeders. The MILP formulation relies on decomposition to rectangular coordinate for complex variables and its accuracy degrades. The exhaustive search, by operating directly on complex values, achieves lower voltage errors and allows higher reduction levels. In early iterations, the error surface is typically flat, meaning many candidate reductions yield negligible voltage error. For example, in the first iteration of Opti-KRON for network A, there are 32 distinct reductions whose effects on the objective function are nearly indistinguishable ($10^{-6}$). This flatness allows the optimizer to select among multiple feasible reductions with similar accuracy, which alters the adjacency matrix across iterations and can lead to different reduced networks.
@@ VI-B Accuracy and generalization
In this subsection, the accuracy and robustness of the reduced networks of the full Vermont feeders. Figure 6 illustrates the trade-off between network reduction and voltage accuracy for the full A and B feeders.
@@
For meshed reduced networks, each point represents the converged reduced network from the exhaustive search for a voltage-deviation limit $\bar{E}\in[0.0001,0.01]$ p.u. For example, $\bar{E}=0.001$ p.u. yields 65% and 81% for reduction for Networks A and B, respectively. Both networks demonstrate a monotonic decrease in accuracy with increasing levels of reduction.
The difference in reduction between two networks for the same error threshold arises from the distinct network structures and loading conditions. These differences lead to variations in nodal voltages and different sensitivities to reduction.
@@ VII Conclusion and Future Work
This work extended the network-reduction framework previously developed for single-phase feeders to unbalanced three-phase distribution systems. The proposed formulation enables accurate aggregation of large feeders while maintaining the voltage profile of the original networks across all phases. By adopting a direct exhaustive-search approach and leveraging parallel execution on GPUs, the method achieves high reduction levels with negligible voltage deviation. We applied the proposed framework into two real-word unbalance three-phase networks and demonstrated the scalability and practicality of the technique.
Future work will generalize the Opti-KRON framework to transmission systems, with emphasis on preserving line flow accuracy between the reduced and original networks. We are also interested in investigating how optimization and control tasks—such as optimal power flow and hosting capacity assessment—performed on reduced models, can be consistently lifted back to the full network.

### v2

@@ VI Simulation and Results
To evaluate the performance of the proposed framework, simulations are carried out on two realistic three-phase unbalanced feeders from Vermont, USA: feeder A with 8,381 nodes and feeder B with 5,591 nodes. To investigate performance across different network scales, multiple sub-feeders of varying sizes (100, 300, 500, and 1000 nodes) are extracted from these systems. For each network, two representative operating scenarios are considered. The two scenarios correspond to the highest and lowest loading conditions observed over a week of historical hourly data, i.e., 168 loading profiles. All computations were performed on the Vermont Advanced Computing Center (VACC) at the University of Vermont. The simulation codes used in this study were implemented in Julia 1.11.5. To solve the optimization problems, JuMP v1.29.1 [undefp] and the Gurobi Optimizer 12.0 were employed, with the optimality gap set to 0%. The GPU implementation was carried out using the CUDA package [undefq].
@@
Across each test feeder, the reduction levels and voltage errors ($0.001$ p.u.) are almost identical, which emphasizes that increasing $q$ provides no meaningful gain in accuracy or reduction capability. On the other hand, small variations in SMICE suggest that larger $q$ values may slightly increase intra-cluster error due to reduced exploration of feasible assignments in each iteration. The remainder of the analysis therefore focuses exclusively on $q=1$, which provides minimal computational burden.
Figure 5 compares the MILP-based and exhaustive search formulations on the 100-node sub-feeders. The MILP formulation relies on decomposition into rectangular coordinates for complex variables, which degrades the accuracy. The exhaustive search, by operating directly on complex values, achieves lower voltage errors and allows higher reduction levels. In early iterations, the error surface is typically flat, meaning many candidate reductions yield negligible voltage error. For example, in the first iteration of Opti-KRON for network A, there are 32 distinct reductions whose effects on the objective function are nearly indistinguishable ($10^{-6}$). This flatness allows the optimizer to select among multiple feasible reductions with similar accuracy, which alters the adjacency matrix across iterations and can lead to different reductions. The MILP formulation serves as a methodological baseline. However, its combinatorial complexity limits scalability to large three-phase feeders. For larger feeders, we therefore rely on the exhaustive-search approach, which scales significantly better in practice.
@@ VI-B Accuracy and generalization
In this subsection, the accuracy and robustness of the reduced networks of the full Vermont feeders are investigated. Figure 6 illustrates the trade-off between network reduction and voltage accuracy for the full A and B feeders.
@@
For meshed reduced networks, each point represents the converged reduced network from the exhaustive search for a voltage-deviation limit $\bar{E}\in[0.0001,0.01]$ p.u. For example, $\bar{E}=0.001$ p.u. yields 65% and 81% reduction for networks A and B, respectively. Both networks demonstrate a monotonic decrease in accuracy with increasing levels of reduction.
The difference in reduction between the two networks for the same error threshold arises from the distinct network structures and loading conditions. These differences lead to variations in nodal voltages and different sensitivities to reduction.
@@ VII Conclusion and Future Work
This work extended the network-reduction framework previously developed for single-phase feeders to unbalanced three-phase distribution systems. The proposed formulation enables accurate aggregation of large feeders while maintaining the voltage profile of the original networks across all phases. By adopting a direct exhaustive-search approach and leveraging parallel execution on GPUs, the method achieves high reduction levels with negligible voltage deviation. We applied the proposed framework to two real-world unbalanced three-phase networks and demonstrated the scalability and practicality of the technique.
Future work will generalize the Opti-KRON framework to transmission systems, with emphasis on preserving line flow accuracy between the reduced and original networks. We are also interested in investigating how optimization and control tasks—such as optimal power flow and hosting capacity assessment—performed on reduced models can be consistently lifted back to the full network.